OpenStudy (anonymous):

Can anyone help me? For the function f(x)=-(3.4)^(x)+8 find f(0), f(-6), and f(4) @Algebra

OpenStudy (stapeleliz):

Just plug the given numbers into the given formula. To find f(0), plug "0" in for "x", and simplify. To find f(-6), plug "-6" in for "x", and simplify. Etc.

OpenStudy (asnaseer):

think of f(x) as a "number factory". it is a factory that, given some number, will return another number using the formulae given. so f(0) means "what number will the number factory f(x) return if I give it a value of x=0" does that give you enough information to work out the answers?

OpenStudy (anonymous):

Yes, I think so. Does that mean the equation will look like x=-(3.4)^(0)+8 or do I plug 0 in for x at the beginning too?

OpenStudy (asnaseer):

almost, the correct form would be:\[f(0)=3.4^0+8=1+8=9\]you can do the same for the other cases.

OpenStudy (anonymous):

Thank you!

OpenStudy (asnaseer):

sorry I made a small error and missed the minus sign, it should be:\[f(0)=-3.4^0+8=-1+8=7\]assuming the original function is:\[f(x)=-3.4^x+8\]

OpenStudy (anonymous):

yes, that is correct

OpenStudy (asnaseer):

ok, hopefully you can do the rest. let me know if you need any help.